OpenStudy (anonymous):

Limit question:

OpenStudy (anonymous):

\[\Large \lim_{x \rightarrow 0} \frac{1}{x^{3}} \int\limits_{0}^{x} \ln(1-t^{2})dt\]

OpenStudy (anonymous):

i tried evaluating the limit hoping that something in the evaluation would cancel out the x^3 in the denominator but i had no such luck...any advice?

OpenStudy (anonymous):

@amistre64 @experimentX

OpenStudy (tkhunny):

Looks like an Indeterminant Form, to me. Type 0/0.

OpenStudy (sirm3d):

yes, 0/0. just apply lhopital's rule

OpenStudy (amistre64):

\[\Large \lim_{x \to 0} \frac{1}{x^{3}} \int\limits_{0}^{x} \ln(1-t^{2})dt\] \[\Large \frac{\frac d{dx}\int\limits_{0}^{x} \ln(1-t^{2})dt}{\frac d{dx} x^3}\] \[\Large \frac{ln(1-x^2)}{3x^2}\] \[\Large \frac{\frac d{dx}ln(1-x^2)}{\frac d{dx}3x^2}\] \[\Large \frac{\frac{-2x}{1-x^2}}{6x}\] \[\Large \frac{-1}{3(1-x^2)}\] now as x goes to zero?

OpenStudy (amistre64):

the key is to recognize the LHop; and also what it means to take the derivative of an integration \[\frac d{dx}~\int_{a(x)}^{b(x)}{f(t)~dt}\to ~\frac d{dx}~[F(b(x))-F(a(x))]\] \[\frac d{dx}~[F(b(x))-F(a(x))]\to~f(b(x))b'-f(a(x))a'\] in this case; f(t) = ln(1-t^2) therefore \[\frac d{dx}~[F(b(x))-F(a(x))]\to~ln(1-x^2)-ln(1-0^2)\]

OpenStudy (amistre64):

pfft, and i spose 0' = 0 so that would also just cancel the f(a) bits regardless :)

OpenStudy (anonymous):

Thanks, that's very helpful. What would your final answer be for the question?

OpenStudy (amistre64):

uhhh .... as long as i didnt mess up along the way; im sticking with my original solution \[\Large \lim_{x\to0} ~\frac{-1}{3(1-x^2)}\]

OpenStudy (anonymous):

the solution i have is 3/10

OpenStudy (amistre64):

the solution that you came up with? or is it a solution from a key?